# CLIMB-3D: Continual Learning for Imbalanced 3D Instance Segmentation

### v2

@@ Abstract
While 3D instance segmentation (3DIS) has advanced significantly, existing methods typically assume that all object classes are known in advance and are uniformly distributed. However, this assumption is unrealistic in dynamic, real-world environments where new classes emerge gradually and exhibit natural imbalance. Although some approaches have addressed class emergence, they often overlook class imbalance, resulting in suboptimal performance – particularly on rare categories. To tackle this challenge, we propose CLIMB-3D, a unified framework for CLass-incremental Imbalance-aware 3DIS. Building upon established exemplar replay (ER) strategies, we show that ER alone is insufficient to achieve robust performance under constrained memory conditions. To mitigate this, we introduce a novel pseudo-label generator (PLG) that extends supervision to previously learned categories by leveraging predictions from a frozen prior model. Despite its promise, PLG tends to bias towards frequent classes. Therefore, we propose a class-balanced re-weighting (CBR) scheme, that estimates object frequencies from pseudo-labels and dynamically adjusts training bias – without requiring access to past data. We design and evaluate three incremental scenarios for 3DIS on the challenging ScanNet200 dataset, and additionally on semantic segmentation on ScanNetV2. Our approach achieves state-of-the-art results, surpassing prior work by up to 16.76% mAP for instance segmentation and approximately 30% mIoU for semantic segmentation, demonstrating strong generalization across both frequent and rare classes.
@@ 1 Introduction
Figure 1: Overview of the CI-3DIS setting: new classes are introduced incrementally as tasks progress. After each task, the model must recognize both previously learned and newly added classes. For instance, at Task 2, classes like Pillow, Coffee Table, and Sofa Chair are introduced, while the model retains the ability to detect earlier classes such as Table, Chair, and Couch.
3D instance segmentation (3DIS) is an essential task in computer vision with widespread applications in graphics, robotics, and augmented reality [boudjoghra20243d, lai2022stratified]. It involves identifying and segmenting individual objects in 3D space, enabling precise object boundaries and class labels that enhance scene understanding and perception in dynamic environments.
@@
In recent years, a variety of methods for 3DIS have been proposed, including top-down [wang2018sgpn, jia2021scaling, zhang2021point], bottom-up [hou20193dsis, yang2019learning], and transformer-based approaches [Schult23ICRA]. These methods have demonstrated strong performance under the traditional setting, which assumes that all object classes are available and well-balanced during training. However, these assumptions limit their applicability in real world settings, where new categories gradually emerge over time, often exhibiting natural class imbalance.
@@
These limitations in current 3DIS methods underscore the need for a Class-Incremental Learning (CIL) [rebuffi2017icarl, de2021continual] framework capable of learning new categories as they appear, while maintaining accuracy on previously learned classes – especially rare ones, which are more susceptible to without catastrophic forgetting [mccloskey1989catastrophic]. Unfortunately, most existing research in CIL has focused on the 2D domain for image classification [rebuffi2017icarl, li2017learning, aljundi2018memory, serra2018overcoming], with limited extensions to object detection [liu2023continual, joseph2021towards, shmelkov2017incremental] and semantic segmentation [cermelli2020modeling, cermelli2022incremental, douillard2021plop]. Extensions to 3D point clouds remain limited, with most work focusing on object-level classification rather than scene-level segmentation [dong2021i3dol, liu2021l3doc, chowdhury2021learning]. Recent works on scene-level class-incremental 3DIS (CI-3DIS) [boudjoghra20243d] and class-incremental semantic segmentation (CI-3DSS) [Yang_2023_CVPR] show promise, but typically rely on large exemplar buffers [rolnick2019experience] Moreover, they typically overlook class imbalance, limiting real-world applicability.
@@
To address this, we propose CLIMB-3D, a unified framework for CLass-incremental IMBalance-aware 3DIS that jointly tackles catastrophic forgetting and class imbalance. The framework begins with Exemplar Replay (ER), storing a small number of samples from past classes for later replay. However, this alone does not yield promising results under the strict memory constraints needed CI-3DIS. To address this, we introduce a Pseudo-Label Generator (PLG), which uses a frozen previous task model to generate supervision for earlier classes. However, we observed that, PLG tends to favour frequent classes while ignoring under-represented ones. To mitigate this, we propose a Class-Balanced Re-weighting (CBR), which estimates class frequencies from the pseudo-labels to enhance the learning of rare classes without accessing past data. Together, these components form a practical and effective solution for real-world 3DIS.
@@
To evaluate CLIMB-3D in a realistic incremental setup, we design three benchmark scenarios using the ScanNet200 [rozenberszki2022language] dataset, which contains 200 classes with a natural long-tail distribution. These scenarios simulate real-world conditions where new classes gradually emerge, following inherently imbalanced distributions. Specifically, new classes appear in groups based on one of three criteria: (A) object frequency, (B) semantic similarity, or (C) random grouping without a specific pattern. Our experiments on these scenarios for incremental 3DIS show that CLIMB-3D effectively mitigates forgetting of previous tasks while improving performance across both frequent and rare classes, delivering consistent gains over prior methods.
@@
In summary, our contributions are threefold: (i) We introduce a novel problem setting of imbalanced class-incremental 3DIS, along with a simple and effective method to mitigate catastrophic forgetting and ensure balanced learning; (ii) We establish three benchmark scenarios that realistically model continuous object emergence with natural class imbalance; and (iii) We validate our framework’s effectiveness, achieving SOTA results with improvements up to 16.76% mAP over existing baselines [^1] Code will be available at: https://github.com/vgthengane/CLIMB3D .
@@ 2 Related Work
This section reviews the current literature on 3D instance segmentation and incremental learning methods, including the limited work addressing incremental learning for 3D scene-level tasks.
2.1 3D Instance Segmentation
Various approaches have been proposed for 3D instance segmentation. One common approach adopts a bottom-up pipeline, in which an embedding in the latent space is learned to facilitate the clustering of object points [wang2018sgpn, lahoud20193d, jiang2020pointgroup, zhang2021point, chen2021hierarchical, han2020occuseg, he2021dyco3d, liang2021instance]. These methods are also known as grouping-based or clustering-based methods. Other methods use a top-down approach, also known as proposal-based methods, where 3D bounding boxes are first detected, then the object region is segmented within the box [yang2019learning, hou20193dsis, liu2020learning, yi2019gspn, engelmann20203d]. Recently, the transformer architecture [vaswani2017attention] has also been used for the task of 3D instance segmentation [Schult23ICRA, sun2022superpoint], motivated by work in 2D [cheng2022masked, cheng2021per]. While these methods propose various models for improving the quality of the object segments, they rely on the availability of annotations for all object categories. On the other hand, we target learning in a progressive manner, in which new semantic annotation is provided and past data is inaccessible.
In order to reduce the annotation cost for 3D instance segmentation, various methods propose weakly supervised alternatives to methods that use dense annotations [xie2020pointcontrast, hou2021exploring, chibane2022box2mask]. While these methods improve the ability to learn from a small set of annotated examples, they rely on a fixed set of semantic labels, so they are prone to catastrophic forgetting in an incremental setting.
2.2 Incremental Learning
Incremental, lifelong, or continual learning methods aim to train a machine learning model sequentially to avoid “catastrophic forgetting” which is caused by training the model on a set of data and later training on another set of data. There are several methods have been proposed for this paradigm, these methods can be divided into three categories: (i) Model Regularization [kirkpatrick2017overcoming, aljundi2018memory, li2017learning] methods limit the plasticity of model parameters to avoid catastrophic forgetting of previous tasks. These methods include weight regularization such as EWC [serra2018overcoming] and function regularization such as knowledge distillation [hinton2015distilling]. (ii) Exemplar replay approaches either create a subset of the past task data or generate samples using generative models to avoid privacy concerns and save those in memory to replay while learning new tasks. This method is effective in more challenging settings and datasets [rebuffi2017icarl, kamra2017deep, buzzega2020dark, cha2021co2l]. (iii) Dynamic network expansion-based method learns a new task by either dynamically expanding the model [rusu2016progressive, li2019learn, zhao2022deep] or by creating a subset of the model [zhao2022deep, wang2020learn, ke2020continual, rajasegaran2019random] to learn to cater for a new task.
Recent approaches to 3D class-incremental segmentation, such as [Yang_2023_CVPR] and [su2024balanced], have made some initial contributions. However, these methods often fall short in performance as they do not leverage state-of-the-art 3D segmentation models and are primarily focused on semantic segmentation, while our work emphasizes object-level instance segmentation. Kontogianni et al. [kontogianni2024continual] propose a general online continual learning framework and evaluate it on 3D dataset segmentation. Similarly, [boudjoghra20243d] addresses the open-world 3D incremental learning problem but relies heavily on an extensive memory buffer. In contrast, our work introduces a dedicated continual learning framework for 3D instance segmentation, with a focus on effective knowledge transfer from previous tasks, while also accounting for the challenges posed by infrequent class occurrences.
@@ 3 Methodology
Figure 2: Overview of CLIMB-3D for CI-3DIS. The model incrementally learns new classes in sequential phases using ER to mitigate forgetting, a PLG for supervision on old classes, and CBR to handle class imbalance.
@@ 3.1 Problem Formulation
3DIS. The objective of this task is to detect and segment each object instance within a given point cloud. Formally, the training dataset is denoted by $\mathcal{D}=\{(\mathbf{P}_{i},\mathbf{Y}_{i})\}_{i=1}^{N}$, where $N$ is the number of training samples. Each input $\mathbf{P}_{i}\in\mathbb{R}^{M\times 6}$ is a colored point cloud consisting of $M$ points, where each point is represented by its 3D coordinates and RGB values. The corresponding annotation $\mathbf{Y}_{i}=\{(m_{i,j},c_{i,j})\}_{j=1}^{J_{i}}$ contains $J_{i}$ object instances, where $m_{i,j}\in\{0,1\}^{M}$ is a binary mask indicating the points belonging to the $j$-th instance, and $c_{i,j}\in\mathcal{C}=\{1,\dots,C\}$ is the semantic class label of that instance. Given $\mathbf{P}_{i}$, the model $\Phi$ predicts instance-level outputs $\hat{\mathbf{Y}}_{i}=\{(\hat{m}_{i,j},\hat{c}_{i,j})\}_{j=1}^{\hat{J}_{i}}$, where $\hat{m}_{i,j}$ and $\hat{c}_{i,j}$ denote the predicted mask and category label for the $j$-th instance, respectively. The number of predicted instances $\hat{J}_{i}$ varies depending on the model’s inference. The model is optimized using below objective:
$\mathcal{L}_{\text{3DIS}}(\mathcal{D};\Phi)$ $=\frac{1}{|\mathcal{D}|}\sum_{(\mathbf{P},\mathbf{Y})\in\mathcal{D}}\frac{1}{|\mathbf{Y}|}\sum_{(m_{j},c_{j})\in\mathbf{Y}}$
$\quad\left(\mathcal{L}_{\text{mask}}(m_{j},\hat{m}_{j})+\lambda\cdot\mathcal{L}_{\text{cls}}(c_{j},\hat{c}_{j})\right)$ (1)
@@
CI-3DIS. Unlike conventional 3DIS, CI-3DIS involves sequential learning, where object categories are introduced incrementally over $T$ training tasks. Each task $t$ introduces a disjoint set of classes $\mathcal{C}^{t}$, with $\mathcal{C}=\bigcup_{t=1}^{T}\mathcal{C}^{t}$ and $\mathcal{C}^{t}\cap\mathcal{C}^{t^{\prime}}=\emptyset$ for $t\neq t^{\prime}$. At each task $t$, the model receives a dataset $\mathcal{D}^{t}=\{(\mathbf{P}_{i},\mathbf{Y}_{i}^{t})\}_{i=1}^{N}$, where each coloured point cloud $\mathbf{P}_{i}\in\mathbb{R}^{M\times 6}$ contains $M$ points, and the corresponding annotation $\mathbf{Y}_{i}^{t}=\{(m_{i,j}^{t},c_{i,j}^{t})\}_{j=1}^{J_{i}^{t}}$ includes instance masks $m_{i,j}^{t}\in\{0,1\}^{M}$ and semantic labels $c_{i,j}^{t}\in\mathcal{C}^{t}$. The model $\Phi^{t}$, initialized from $\Phi^{t-1}$, is trained on $\mathcal{D}^{t}$ to predict instance-level outputs $\hat{\mathbf{Y}}_{i}^{t}=\{(\hat{m}_{i,j}^{t},\hat{c}_{i,j}^{t})\}_{j=1}^{\hat{J}_{i}^{t}}$, where each predicted class label $\hat{c}_{i,j}^{t}$ belongs to the cumulative set of all classes observed so far: $\bigcup_{k=1}^{t}\mathcal{C}^{k}$. The optimisation objective for the CI-3DIS setting is defined as:
$\mathcal{L}_{\text{CI-3DIS}}(\mathcal{D}^{t};\Phi^{t})$ $=\frac{1}{|\mathcal{D}^{t}|}\sum_{(\mathbf{P},\mathbf{Y}^{t})\in\mathcal{D}^{t}}\frac{1}{|\mathbf{Y}|}\sum_{(m^{t}_{j},c^{t}_{j})\in\mathbf{Y}^{t}}$
$\quad\left(\mathcal{L}_{\text{mask}}(m^{t}_{j},\hat{m}^{t}_{j})+\lambda\cdot\mathcal{L}_{\text{cls}}(c^{t}_{j},\hat{c}^{t}_{j})\right).$ (2)
The key challenge in this setup is to learn new classes without forgetting those encountered in previous tasks, despite limited supervision and the absence of past labels.

### v3

@@ Abstract
While 3D instance segmentation (3DIS) has advanced significantly, most existing methods assume that all object classes are known in advance and uniformly distributed. However, this assumption is unrealistic in dynamic, real-world environments where new classes emerge gradually and exhibit natural imbalance. Although some approaches address the emergence of new classes, they often overlook class imbalance, which leads to suboptimal performance, particularly on rare categories. To tackle this, we propose CLIMB-3D, a unified framework for CLass-incremental Imbalance-aware 3DIS. Building upon established exemplar replay (ER) strategies, we show that ER alone is insufficient to achieve robust performance under memory constraints. To mitigate this, we introduce a novel pseudo-label generator (PLG) that extends supervision to previously learned categories by leveraging predictions from a frozen model trained on prior tasks. Despite its promise, PLG tends to be biased towards frequent classes. Therefore, we propose a class-balanced re-weighting (CBR) scheme that estimates object frequencies from pseudo-labels and dynamically adjusts training bias, without requiring access to past data. We design and evaluate three incremental scenarios for 3DIS on the challenging ScanNet200 dataset and additionally validate our method for semantic segmentation on ScanNetV2. Our approach achieves state-of-the-art results, surpassing prior work by up to 16.76% mAP for instance segmentation and approximately 30% mIoU for semantic segmentation, demonstrating strong generalisation across both frequent and rare classes. Code is available at: https://github.com/vgthengane/CLIMB3D.
@@ 1 Introduction
Figure 1: Overview of the CI-3DIS setting. New object categories are introduced incrementally with each task. After every phase, the model must recognise both newly added and previously learned classes. For instance, in Task 2, categories such as Pillow, Coffee Table, and Sofa Chair are introduced, while the model is expected to retain recognition of earlier classes like Table, Chair, and Couch.
@@
3D instance segmentation (3DIS) is a fundamental task in computer vision that involves identifying precise object boundaries and class labels in 3D space, with broad applications in graphics, robotics, and augmented reality [boudjoghra20243d, lai2022stratified]. Traditional 3DIS methods, including top-down [wang2018sgpn, jia2021scaling, zhang2021point], bottom-up [hou20193dsis, yang2019learning], and transformer-based approaches [Schult23ICRA], perform well under the assumption that all classes are available and balanced during training. However, this assumption does not hold in real-world environments, where new object classes emerge over time and exhibit natural imbalance.
@@
This gap motivates the need for Class-Incremental Learning (CIL) [rebuffi2017icarl, de2021continual], which supports learning new categories while retaining knowledge of previous ones [mccloskey1989catastrophic]. While CIL has seen success in 2D image tasks [rebuffi2017icarl, li2017learning, aljundi2018memory, serra2018overcoming], extensions to 3D point clouds remain limited, often focusing on object-level classification [dong2021i3dol, liu2021l3doc, chowdhury2021learning]. Recent works on scene-level class-incremental 3DIS (CI-3DIS) [boudjoghra20243d] and class-incremental semantic segmentation (CI-3DSS) [Yang_2023_CVPR] show promise, but depend heavily on large exemplar [rolnick2019experience] and overlook class imbalance, limiting their practicality.
@@
To address this, we propose CLIMB-3D, a unified framework for CLass-incremental IMBalance-aware 3DIS that jointly tackles catastrophic forgetting and class imbalance. The framework begins with Exemplar Replay (ER), storing a small number of samples from past classes for later replay. However, this alone does not yield promising results under the strict memory constraints required in CI-3DIS. To address this, we introduce a Pseudo-Label Generator (PLG), which uses a frozen model from previous task to generate supervision for earlier classes. However, we observed that PLG tends to favour frequent classes while ignoring under-represented ones. To mitigate this, we propose a Class-Balanced Re-weighting (CBR), which estimates class frequencies from the pseudo-labels to enhance the learning of rare classes without accessing past data. Together, these components form a practical and effective solution for real-world 3DIS.
@@
To evaluate CLIMB-3D, we designed three benchmark scenarios for CI-3DIS based on the ScanNet200 dataset [rozenberszki2022language]. These scenarios simulate incremental learning setup under natural class imbalance, where new classes emerge based on (A) object frequency, (B) semantic similarity, or (C) random grouping. Additionally, for comparison with existing methods, we evaluate our approach on CI-3DSS using the ScanNetV2 dataset [dai2017scannet]. Experimental results show that our method significantly reduces forgetting and improves performance across both frequent and rare categories, outperforming previous methods in both settings.  Fig.˜1 illustrates the CI-3DIS setup.
@@
In summary, our contributions are: (i) a novel problem setting of imbalanced class-incremental 3DIS with an effective method to balance learning and mitigate forgetting; (ii) three benchmarks modelling continual object emergence with natural imbalance; and (iii) strong experimental results, achieving up to 16.76% mAP improvement over baselines.
@@ 2 Related Work
3D Instance Segmentation. Various methods have been proposed for 3DIS. Grouping-based approaches adopt a bottom-up pipeline that learns latent embeddings for point clustering [wang2018sgpn, lahoud20193d, jiang2020pointgroup, zhang2021point, chen2021hierarchical, han2020occuseg, he2021dyco3d, liang2021instance]. In contrast, proposal-based methods follow a top-down strategy by detecting 3D boxes and segmenting objects within them [yang2019learning, hou20193dsis, liu2020learning, yi2019gspn, engelmann20203d]. Recently, transformer-based models [vaswani2017attention] have also been applied to 3DIS and 3DSS [Schult23ICRA, sun2022superpoint], inspired by advances in 2D vision [cheng2022masked, cheng2021per]. However, these approaches require full annotations for all classes and are not designed for progressive learning, where only new class annotations are available and previous data is inaccessible. Another line of work aims to reduce annotation costs by proposing weakly supervised 3DIS methods based on sparse cues [xie2020pointcontrast, hou2021exploring, chibane2022box2mask]. While effective with limited annotations, these methods assume a fixed set of classes and are susceptible to catastrophic forgetting in incremental settings.
Incremental Learning. Continual learning involves training models sequentially to mitigate catastrophic forgetting. One common strategy is model regularisation, which constrains parameter updates using techniques such as elastic weight consolidation or knowledge distillation [kirkpatrick2017overcoming, aljundi2018memory, li2017learning, serra2018overcoming, hinton2015distilling]. Another approach is exemplar replay, where past data is stored or generated to rehearse older tasks while learning new ones [rebuffi2017icarl, kamra2017deep, buzzega2020dark, cha2021co2l]. A third direction involves dynamically expanding the model architecture or using modular sub-networks to accommodate new tasks without interfering with old ones [rusu2016progressive, li2019learn, zhao2022deep, wang2020learn, ke2020continual, rajasegaran2019random]. Recent efforts in class-incremental 3D segmentation [Yang_2023_CVPR, su2024balanced] have shown early promise but often rely on basic architectures and focus primarily on semantic segmentation. Other works [kontogianni2024continual, boudjoghra20243d] explore continual learning in 3D settings, though they often depend on large memory buffers. In contrast, our work introduces a tailored framework for 3D instance segmentation that effectively transfers knowledge across tasks, even under class imbalance.
Long-tailed Recognition. Imbalanced class distributions lead to poor recognition of rare (long-tail) classes. To address this, re-sampling methods balance the data via input [chawla2002smote, gupta2019lvis, park2022majority, shen2016relay, mahajan2018exploring] or feature space [ando2017deep, chu2020feature, li2021metasaug], avoiding naïve under-/over-sampling. Loss re-weighting offers another direction, using class-based [khan2017cost, he2022relieving, cao2019learning, khan2019striking, cui2019class, wang2021adaptive] or per-example [lin2017focal, ren2018learning, shu2019meta] adjustments to ensure fair contribution. Parameter regularisation improves generalisation through weight constraints [alshammari2022long], albeit requiring careful tuning. Other approaches leverage transfer learning [yin2019feature, zhou2020bbn], self-supervision [yang2020rethinking, li2021self], or contrastive learning [khosla2020supervised, li2021self, zhu2022balanced] to improve rare-class representations. Long-tailed recognition is well-studied in 2D with large-scale datasets [cui2019class, van2018inaturalist, liu2019large]; in 3D, ScanNet200 [rozenberszki2022language, dai2017scannet] enables similar efforts. Prior 3D work focused on re-weighting, re-sampling, and transfer learning [rozenberszki2022language], while regularisation remains underexplored. CeCo [zhong2023understanding] balances class centres via auxiliary loss but overlooks sampling and augmentation. Lahoud et al. [lahoud2024long] propose adaptive classifier regularisation for 3D segmentation, outperforming prior approaches without threshold tuning. However, neither method considers incremental learning. In contrast, we jointly address long-tail and continual 3D semantic segmentation.
@@ 3 Methodology
Figure 2: Overview of CLIMB-3D for CI-3DIS. The model incrementally learns new classes across sequential phases. During task $t$, point clouds $\mathbf{P}$ and their corresponding labels $\mathbf{Y}^{t}$ are sampled from a combination of the current training dataset $D^{t}$ and Exemplar Replay (ER), which maintains a small memory of past examples. These are then passed to the current model $\Phi^{t}$ to produce predicted labels $\mathbf{Y}^{t}$. The Pseudo-Label Generator (PLG) selects the top-$K$ predictions from the previous model $\Phi^{t-1}$. These pseudo-labels are then weighted based on class frequency $f(c)$ using Class-Balanced Re-weighting (CBR), and the top-$K$ re-weighted labels are selected to form the balanced pseudo-label set $\bar{\mathbf{Y}}^{t}$. This pseudo-label set is then concatenated with the ground-truth labels to form a final augmented supervision set $\overline{\mathbf{Y}}^{t}$ for task $t$, which is used to optimise the model $\Phi^{t}$ using Eq.˜3.
@@ 3.1 Problem Formulation
3DIS. The objective of this task is to detect and segment each object instance within a point cloud. Formally, the training dataset is denoted as $\mathcal{D}=\{(\mathbf{P}_{i},\mathbf{Y}_{i})\}_{i=1}^{N}$, where $N$ is the number of training samples. Each input $\mathbf{P}_{i}\in\mathbb{R}^{M\times 6}$ is a coloured point cloud consisting of $M$ points, where each point is represented by its 3D coordinates and RGB values. The corresponding annotation $\mathbf{Y}_{i}=\{(m_{i,j},c_{i,j})\}_{j=1}^{J_{i}}$ contains $J_{i}$ object instances, where $m_{i,j}\in\{0,1\}^{M}$ is a binary mask indicating which points belong to the $j$-th instance, and $c_{i,j}\in\mathcal{C}=\{1,\dots,C\}$ is the semantic class label of that instance. Given $\mathbf{P}_{i}$, the model $\Phi$ predicts instance-level outputs $\hat{\mathbf{Y}}_{i}=\{(\hat{m}_{i,j},\hat{c}_{i,j})\}_{j=1}^{\hat{J}_{i}}$, where $\hat{m}_{i,j}$ and $\hat{c}_{i,j}$ denote the predicted mask and category label for the $j$-th instance, respectively. The number of predicted instances $\hat{J}_{i}$ varies depending on the model’s inference. The model is optimised using the following objective:
$\mathcal{L}_{\text{3DIS}}(\mathcal{D};\Phi)=\frac{1}{|\mathcal{D}|}\sum_{(\mathbf{P},\mathbf{Y})\in\mathcal{D}}\frac{1}{|\mathbf{Y}|}\sum_{(m_{j},c_{j})\in\mathbf{Y}}\left(\mathcal{L}_{\text{mask}}(m_{j},\hat{m}_{j})+\lambda\cdot\mathcal{L}_{\text{cls}}(c_{j},\hat{c}_{j})\right),$ (1)
@@
CI-3DIS. Unlike conventional 3DIS, CI-3DIS involves sequential learning, where object categories are introduced incrementally over $T$ training tasks. Each task $t$ introduces a disjoint set of classes $\mathcal{C}^{t}$, with $\mathcal{C}=\bigcup_{t=1}^{T}\mathcal{C}^{t}$ and $\mathcal{C}^{t}\cap\mathcal{C}^{t^{\prime}}=\emptyset$ for $t\neq t^{\prime}$. During each task $t$, the model receives a dataset $\mathcal{D}^{t}=\{(\mathbf{P}_{i},\mathbf{Y}_{i}^{t})\}_{i=1}^{N}$, where each coloured point cloud $\mathbf{P}_{i}\in\mathbb{R}^{M\times 6}$ contains $M$ points, and the corresponding annotation $\mathbf{Y}_{i}^{t}=\{(m_{i,j}^{t},c_{i,j}^{t})\}_{j=1}^{J_{i}^{t}}$ includes instance masks $m_{i,j}^{t}\in\{0,1\}^{M}$ and semantic class labels $c_{i,j}^{t}\in\mathcal{C}^{t}$. The model $\Phi^{t}$, initialized from $\Phi^{t-1}$, is trained on $\mathcal{D}^{t}$ to predict instance-level outputs $\hat{\mathbf{Y}}_{i}^{t}=\{(\hat{m}_{i,j}^{t},\hat{c}_{i,j}^{t})\}_{j=1}^{\hat{J}_{i}^{t}}$, where each predicted semantic label $\hat{c}_{i,j}^{t}$ belongs to the cumulative set of all classes observed so far: $\bigcup_{k=1}^{t}\mathcal{C}^{k}$. The key challenge in this setup is to learn new classes without forgetting those encountered in previous tasks, despite limited supervision and the absence of past labels.